You have a 3.00-liter container filled with N₂ at 25°C and 1.75 atm pressure connected to a 2.00-liter container filled with Ar at 25°C and 3.05 atm pressure. A stopcock connecting the containers is opened and the gases are allowed to equilibrate between the two containers. What is the density of the final gas mixture? Assume ideal behavior.
Answer #1

Ar P₂=3.05 N2 P, = 1.15 atm Vi=3L T= 25+273=298k n= PV V2=2L T = 298k RT = 0. 249 = 0.214 n=1.75*3 0.0821x298 n=3.05*2 0.0821